OpenStudy (anonymous):

A function f is defined in such a way that f(1)=2, and for each positive integer n>1, f(1)+f(2)+f(3)+...+f(n)=n^2f(n), evaluate the exact value of f(1000).

OpenStudy (kinggeorge):

I haven't proved anything, but I have noticed a striking pattern. For values above \(n=2\), it appears that \[f(n)={f(1)\over T_n}\]where \(T_n\) is the \(n\)-th triangular number. This would give \[\large f(1000)={2 \over {1000^2-1000 \over 2}}={1\over 999000}\]

OpenStudy (asnaseer):

You can rearrange to get this relationship:\[\begin{align} f(1)+f(2)+...+f(n)&=n^2f(n)\\ f(1)+f(2)+...+f(n-1)&=n^2f(n)-f(n)=(n^2-1)f(n)\\ (n-1)^2f(n-1)&=(n^2-1)f(n)\\ (n-1)f(n-1)&=(n+1)f(n)\\ f(n)&=\frac{n-1}{n+1}f(n-1) \end{align}\]

OpenStudy (asnaseer):

which then leads to....

OpenStudy (asnaseer):

\[\begin{align} f(n)&=\frac{n-1}{n+1}f(n-1)\\ &=\frac{n-1}{n+1}\frac{n-2}{n}f(n-2)\\ &=\frac{n-1}{n+1}\frac{n-2}{n}\frac{n-3}{n-1}f(n-3)\\ &...\\ &=\frac{n-1}{n+1}\frac{n-2}{n}\frac{n-3}{n-1}...\frac{1}{3}f(1)\\ &=\frac{2(n-1)!}{(n+1)!}f(1)\\ &=\frac{2}{n(n+1)}f(1) \end{align}\]

OpenStudy (asnaseer):

from which you should be able to calculate f(1000)

OpenStudy (kinggeorge):

So the pattern I noticed was correct!

OpenStudy (asnaseer):

but I get:\[f(1000)=\frac{2}{1000*1001}*2=\frac{1}{250250}\]

OpenStudy (asnaseer):

I believe your pattern results in:\[f(1000)={2 \over {1000^2-1000 \over 2}}={1\over 249750}\]

OpenStudy (kinggeorge):

I just didn't use the formula for triangular numbers >.< It should be \[\large {2 \over {1000^2+1000 \over 2}}\]And I calculated it wrong.

OpenStudy (asnaseer):

ah yes! - you're right - I should have spotted that as well :)